# Search for $C$-even states decaying to $D_{s}^{\pm}D_{s}^{*\mp}$ with masses between $4.08$ and $4.32$ $\rm GeV/{\it c}^{2}$

### v1

@@ Abstract
The charmonium system is a good laboratory for studying the nonperturbative behavior of quantum chromodynamics (QCD). In the past two decades, a few resonant structures have been observed that cannot be explained by a simple mesonic ($q\bar{q}$) or baryonic ($qqq$) configuration, which has resulted in great experimental and theoretical interest.
Charmonium states with mass above the open-charm threshold are expected to decay dominantly to open-charm final states  ozione ; ozitwo ; ozithree . In Ref. Higher_Charmonia , the masses and total widths of the $\chi_{c2}(3P)$, $\chi_{c1}(3P)$, and $\eta_{c2}(2D)$ are calculated according to the nonrelativistic potential model or the relativized Godfrey-Isgur model Higher_Charmonia1 . The open-charm strong decay widths are determined assuming harmonic oscillator wave functions and the ${}^{3}{P}_{0}$ decay model. These states are predicted to have sizeable decay widths to the $D_{s}^{\pm}D_{s}^{*\mp}$ final state, as listed in Table 1. None of them have so far been observed in experiment. In an amplitude analysis of the $B^{+}\to J/\psi\phi K^{+}$ decay, a $1^{++}$ state was observed in the $\phi J/\psi$ final state lhca ; lhcb ; lhcc , the average mass of which, $4286^{+8}_{-9}~{}{\rm MeV}/c^{\rm 2}$, is close to that of the predicted $\chi_{c1}(3P)$. In the hybrid meson configuration, several states are predicted by lattice QCD QED1 ; QED2 . The strong decay widths of these hybrid states are computed based on the constituent gluon model constituent_gluon . The $D_{\rm s}^{\pm}D_{\rm s}^{*\mp}$ mode is expected to be the leading decay channel for the $X(4217)$ with $J^{PC}=1^{-+}$ and the $X(4279)$ with $J^{PC}=0^{-+}$, where the numbers in the brackets denote the mass of the hybrid states from theoretical predictions. The resonant parameters of the two states are summarized in Table 1.
Many of the reported structures have masses close to the thresholds of open-charm meson pairs. For example, the $X(3872)$ x3872 , $Z_{c}(3900)^{\pm}$ zc3900a ; zc3900b ; zc3900c , $Z_{c}(4020)^{\pm}$ zc4020a ; zc4020b , and $Z_{cs}(3985)$ zcs3985a ; zcs3985b are near the thresholds of $D^{0}\bar{D}^{*0}$, $D\bar{D}^{*}$, $D^{*}\bar{D}^{*}$, and $D_{s}^{(*)}\bar{D}^{(*)}$, respectively. In the meson molecule picture, Ref. heavy-antiheavy predicted abundant hadron spectra near open-charm thresholds by solving the Bethe-Salpeter equation. In addition, a virtual state, referred to as the $X(4080)$ in this paper, with a mass close to the $D_{s}^{\pm}D_{s}^{*\mp}$ threshold, is expected to decay predominantly to $D_{s}^{\pm}D_{s}^{*\mp}$. Its quantum numbers are $J^{PC}=1^{++}$ and its total width is assumed to be $5~{}\mathrm{MeV}$.
Table 1: Masses ($M$), total widths ($\Gamma$), partial widths ($\Gamma_{D_{s}^{\pm}D_{s}^{*\mp}}$), and quantum numbers of the six $C$-even states.
@@
It has been shown that $C$-even states can be produced in $e^{+}e^{-}$ collisions accompanied by photon emission. For example, the $e^{+}e^{-}\to\gamma X(3872)$ gamma3872 and $\gamma\chi_{c1,2}(1P)$ gammachicj processes have been observed by the BESIII experiment. In this paper, a study of $e^{+}e^{-}\to\gamma D_{s}^{\pm}D_{s}^{*\mp}$ is presented to search for the $C$-even structures mentioned above. Throughout this paper, the charge conjugated mode is implied. This analysis is performed using a data sample collected at center-of-mass energy $\sqrt{s}=(4681.92\pm 0.08\pm 0.29)~{}\mathrm{MeV}$ with an integrated luminosity of $(1667.39\pm 0.21\pm 8.84)~{}\mathrm{pb}^{-1}$ BESIII:lum . The $D_{s}^{*}$ meson is reconstructed using its radiative decay into the $\gamma D_{s}$ final state. The $D_{s}$ meson is reconstructed with the subsequent $D_{s}\to K^{+}K^{-}\pi$ and $K^{0}_{S}K$ decays, with $K^{0}_{S}\to\pi^{+}\pi^{-}$. The two decay modes of the $D_{s}$ are combined into four decay chains (DCs):
@@
Simulated events produced with a geant4-based geant4 Monte Carlo (MC) package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate background contributions. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc ref:kkmca ; ref:kkmcb . The generic MC samples include the production of open-charm processes, the ISR production of vector charmonium(-like) states, and the continuum processes implemented in kkmc ref:kkmca ; ref:kkmcb . The known decay modes are modelled with evtgen ref:evtgen ; ref:evtgena using branching fractions taken from the Particle Data Group (PDG) PDG2018 , and the remaining unknown charmonium decays are modelled with lundcharm ref:lundcharma ; ref:lundcharmb . Final state radiation for charged final state particles is incorporated using the photos package ref:photos .
The signal processes $e^{+}e^{-}\to\gamma X,~{}X\to D_{s}^{\pm}D_{s}^{*\mp}$ and their subsequent decays are modelled with evtgen ref:evtgen ; ref:evtgena . The $D_{s}\to K^{+}K^{-}\pi$ decay is generated according to the amplitude models reported in Refs. ref:da1 ; ref:da2 . The input cross section line shape of $e^{+}e^{-}\to\gamma X$ is assumed to follow those of the $\psi(\rm{4660})$, and is parameterized using a Breit-Wigner function with parameters taken from the PDG PDG2018 . Two groups of samples are generated for the signal processes, Group-I and Group-II.Group-I is produced with the predicted masses and widths of the $X$, while Group-II is with the masses ranging from $4080$ to $4320~{}\mathrm{MeV}/c^{2}$ in steps of $40~{}\mathrm{MeV}/c^{2}$ and widths varying from a set of values: $5,~{}10,~{}50,~{}80,\rm{\rm and}~{}100~{}\mathrm{MeV}$. The quantum numbers $J^{PC}$ are listed in Table 1.
Samples of $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{(*)\pm}D_{s}^{*\mp}$ events are generated to estimate the background contamination. They are modelled using helicity-amplitude models in evtgen ref:evtgen ; ref:evtgena . The production cross section is taken from a previous BESIII measurement  ref:dsstardsstar .
@@
Each $K^{0}_{S}$ candidate is reconstructed from two oppositely charged tracks satisfying $|V_{z}|<20~{}\mathrm{cm}$ and $|\cos\theta|<0.93$. The two charged tracks are assigned as $\pi^{+}\pi^{-}$ without imposing PID criteria. They are constrained to originate from a common vertex and are required to have an invariant mass $(M_{\pi^{+}\pi^{-}})~{}\rm{within~{}the~{}interval}~{}(0.478,~{}0.518)$ GeV/$c^{2}$. The $\chi^{2}$ of the vertex fit is required to be less than $100$. A secondary vertex fit is performed to ensure that the $K^{0}_{S}$ momentum points back to the IP. The decay length of the $K^{0}_{S}$ candidate is required to be greater than twice the vertex resolution.
@@
Photon candidates are identified using isolated showers not associated with charged tracks in the EMC. The deposited energy of each shower must be greater than $25~{}\mathrm{MeV}$ in the barrel region ($|\cos\theta|<0.8$) and greater than $50~{}\mathrm{MeV}$ in the end cap regions ($0.920.86<|\cos\theta|<$ 0.92). The difference between the EMC time and the event start time is required to be within $[0,~{}700]~{}\mathrm{ns}$ to suppress electronic noise and showers unrelated to the event.
@@
The selected $K^{\pm}$, $\pi^{\pm}$, and $K^{0}_{S}$ candidates in the event are combined to reconstruct $D_{s}\to K^{+}K^{-}\pi$ or $D_{s}\to K^{0}_{S}K$. The $D_{s}$ candidates are kept if the invariant masses of the $K^{+}K^{-}\pi$ or $K^{0}_{S}K$ systems, denoted as $M_{D_{s}}$, are within a ±$\pm 150~{}\mathrm{MeV}/c^{2}$ mass window around the nominal $D_{s}$ mass. The $D_{s}^{+}D_{s}^{-}$ pairs are selected requiring that each $K^{\pm}$, $\pi^{\pm}$, and $K^{0}_{S}$ candidate is used in the event reconstruction at most once. If there is more than one combination in an event, the one with $(M_{D_{s}^{+}}+M_{D_{s}^{-}})/2$ closest to the nominal mass of the $D_{s}$ meson PDG2018 is chosen.
@@
In addition to the $D_{s}^{+}D_{s}^{-}$ pair, each signal candidate is required to have at least two extra photons. A six-constraint (6C) kinematic fit is performed, where the four-momentum of the final state particles is constrained to that of the initial $e^{+}e^{-}$ system and the masses of the reconstructed $D_{s}^{+}$ and $D_{s}^{-}$ are constrained to their nominal masses. If more than two photons are found or there is more than one pair of $D_{s}^{+}D_{s}^{-}$ decaying in different DCs in an event, the combination with the minimum $\chi^{2}_{\rm 6C}$ is retained and $\chi^{2}_{\rm 6C}$ is further required to be less than 200, where $\chi^{2}_{\rm 6C}$ is the $\chi^{2}$ from the 6C kinematic fit. To study the background contributions from non-$D_{s}$ processes, a four-constraint kinematic (4C) fit without the two $D_{s}$ mass constraints is applied and the four-momenta of the final state particles are computed using the momenta determined by the 4C kinematic fit, unless otherwise specified.
@@
To suppress background contributions from non-$D_{s}$ processes, a $D_{s}$ mass window of $|M_{D_{s}}-m_{D_{s}}|<21~{}\mathrm{MeV}/c^{2}$, which is about three times the mass resolution, is applied, where $m_{D_{s}}$ is the nominal $D_{s}$ mass. The two selected photon candidates and the two $D_{s}$ mesons can form four combinations of $\gamma D_{s}$. The one with invariant mass ($M_{\gamma D_{s}})$ closest to the nominal mass of the $D_{s}^{*}$ is retained for further analysis. The invariant mass $M^{\prime}_{\gamma D_{s}}=(M_{\gamma D_{s}}-M_{D_{s}}+m_{D_{s}})$ is required to be within the interval $(m_{D_{s}^{*}}\pm 27)~{}\mathrm{MeV}/c^{2}$, where $m_{D_{s}^{*}}$ is the nominal $D_{s}^{*\pm}$ mass taken from the PDG PDG2018 . The variable $M_{\gamma D_{s}}$ formed with the unused photon and $D_{s}$ candidates is required to be outside of this same mass window. The photon not from the $D_{s}^{*}$ decay is labelled $\gamma_{1}$ and its recoil mass spectrum, $M_{\gamma_{1}}^{\rm rec}$, is used to extract the signal yields. $M_{\gamma_{1}}^{\rm rec}$ is calculated by $\sqrt{(E_{\rm cms}-E_{\gamma_{1}})^{2}-(\vec{P}_{\rm cms}-\vec{P}_{\gamma_{1}})^{2}}$, where ($E_{\rm cms},\vec{P}_{\rm cms}$) and ($E_{\gamma_{1}},\vec{P}_{\gamma_{1}}$) are the four-momenta of the initial state BESIII:lum and $\gamma_{1}$, respectively. After imposing all the selection requirements, the dominant background processes are $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{\pm}D_{s}^{*\mp}$ and $(\gamma_{\rm ISR})D_{s}^{*\pm}D_{s}^{*\mp}$. The contributions from other background processes are less than $10\%$ of the total backgrounds. Events with a $D^{(*)}\bar{D}^{(*)}$ pair in the final state are found to be linearly distributed in the recoil mass spectrum of $\gamma_{1}$ by analyzing the generic MC samples.
@@
The $X$ signal yields are extracted using an unbinned maximum likelihood fit to the $M_{\gamma_{1}}^{\rm rec}$ spectrum. The events from the four DCs are combined due to limited statistics. The fit probability density function (PDF) contains four components: signal ($S$), $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{\pm}D_{s}^{*\mp}$ background ($B_{1}$), $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{*\pm}D_{s}^{*\mp}$ background ($B_{2}$), and other remaining backgrounds ($B_{3}$). The function used in the fit is defined as
$f_{\rm sum}=N_{\rm sig}\cdot S+N_{\rm bkg1}\cdot B_{1}+N_{\rm bkg2}\cdot B_{2}+N_{\rm bkg3}\cdot B_{3}.$
Here, $N_{\rm sig}$ and $N_{\rm bkg(1,2,3)}$ represent the numbers of signal events and background events, and they are free parameters in the fit.
The signal PDF is a Breit-Wigner function (BW) with mass $M$ and width $\Gamma$, convoluted with a Double Gaussian and a Gaussian function. The Double Gaussian function, with the parameters $m_{2},m_{1},~{}\sigma_{2},\sigma_{1}$, and $f$, describes the mass resolution and mass shift in the simulation, and the Gaussian function accounts for the mass resolution difference between data and simulation. To get the parameters of the Double Gaussian function, the $M_{\gamma_{1}}^{\rm rec}-M_{\gamma_{1}}^{\rm truth}$ distribution from the MC samples is fitted, where $M_{\gamma_{1}}^{\rm truth}$ is the recoil mass spectrum of the $\gamma_{1}$ distribution at generator level. The parameters of the Gaussian function are obtained by using a control sample, $e^{+}e^{-}\to D_{s}^{\pm}D_{s}^{*\mp}$. The selection of the control sample is similar to that of the signal sample but only one photon is required instead of two. The $D_{s}$ and $D_{s}^{*}$ mass distributions from data are fitted with the line shape from MC samples convoluted with Gaussian functions. The parameters of the Gaussian function representing the difference of the mass resolutions between data and MC simulation are then used as the parameters of the Gaussian function in the extraction of the $X$ signal yield. The $B_{1(2)}$ line shape is determined from the $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{\pm}D_{s}^{*\mp}$ ($e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{*\pm}D_{s}^{*\mp}$) MC simulation and $B_{3}$ is a polynomial function.
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Fig. 1: Fits to the $M_{\gamma_{1}}^{\rm rec}$ spectrum with masses and widths fixed to those of the $\eta_{c2}(2D)$, $\chi_{c1}(3P)$, $\chi_{c2}(3P)$, $X(4080)$, $X(4217)$, and $X(4279)$ as listed in Table 1. The dots with error bars are the data distribution and the blue solid lines are the best fit results.
In addition, upper limits for the signal cross sections as a function of the $X$ masses in the range $[4.08,~{}4.32]~{}\mathrm{GeV}/c^{2}$ are also provided as listed in Table 1. From studies of simulation samples (Group-II), it is found the parameters $m_{2}-m_{1}$, $\sigma_{2}/\sigma_{1}$, and $f$ of the resolution function are very similar for various widths and $J^{PC}$. The parameters are found to be

### v2

@@ Abstract
The charmonium(like) system is a good laboratory for studying the nonperturbative behavior of quantum chromodynamics (QCD). Over the past two decades, many new resonant structures have been discovered that cannot be explained by a simple mesonic ($q\bar{q}$) or baryonic ($qqq$) configurations. This has sparked significant experimental and theoretical interest.
Charmonium states with mass above the open-charm threshold are expected to decay dominantly into open-charm final states  ozione ; ozitwo ; ozithree . In Ref. Higher_Charmonia , the masses and total widths of the $\chi_{c2}(3P)$, $\chi_{c1}(3P)$, and $\eta_{c2}(2D)$ are calculated using either the nonrelativistic potential model or the relativized Godfrey-Isgur model Higher_Charmonia1 . The open-charm strong decay widths are estimated by using harmonic oscillator wave functions and the ${}^{3}{P}_{0}$ decay model. These states are predicted to have sizeable decay widths to the $D_{s}^{\pm}D_{s}^{*\mp}$ final state, as listed in Table 1. However, none of these states have been observed experimentally so far. In an amplitude analysis of the $B^{+}\to J/\psi\phi K^{+}$ decay, a $1^{++}$ state was observed in the $\phi J/\psi$ final state lhca ; lhcb ; lhcc , with an average mass of $4286^{+8}_{-9}~{}{\rm MeV}/c^{\rm 2}$, which is close to the predicted mass of the $\chi_{c1}(3P)$. Its assignment as a $\chi_{c1}(3P)$ state requires further confirmation.
In the hybrid meson ($q\bar{q}g$) configuration, several states are predicted by lattice QCD QED1 ; QED2 . The strong decay widths of these hybrid states are computed based on the constituent gluon model constituent_gluon . The $D_{\rm s}^{\pm}D_{\rm s}^{*\mp}$ mode is expected to be the leading decay channel for the $X(4217)$ with $J^{PC}=1^{-+}$ and the $X(4279)$ with $J^{PC}=0^{-+}$, where the numbers in the brackets denote the theoretical masses of the hybrid states. The resonant parameters of the two states are summarized in Table 1.
The authors of Ref. heavy-antiheavy provide a whole spectrum of heavy-antiheavy hadronic molecules by solving the Bethe-Salpeter equation. A virtual state, referred to as the $X(4080)$ in this paper, with a mass close to the $D_{s}^{\pm}D_{s}^{*\mp}$ threshold, is expected to decay predominantly to $D_{s}^{\pm}D_{s}^{*\mp}$. Its quantum numbers are $J^{PC}=1^{++}$ and its total width is assumed to be $5~{}\mathrm{MeV}$. Many structures from experiments with masses near the thresholds of a pair of open-charm mesons are reported, such as the $X(3872)$ x3872 , $Z_{c}(3900)^{\pm}$ zc3900a ; zc3900b ; zc3900c , $Z_{c}(4020)^{\pm}$ zc4020a ; zc4020b , and $Z_{cs}(3985)$ zcs3985a ; zcs3985b , which are near the thresholds of $D^{0}\bar{D}^{*0}$, $D\bar{D}^{*}$, $D^{*}\bar{D}^{*}$, and $D_{s}^{(*)}\bar{D}^{(*)}$, respectively. Motivated by both experimental observations and theoretical calculations, searching for the structure near the threshold of $D_{s}^{\pm}D_{s}^{*\mp}$ is crucial.
Table 1: Predicted masses ($M$), total widths ($\Gamma$), partial widths ($\Gamma_{D_{s}^{\pm}D_{s}^{*\mp}}$), and quantum numbers of the six $C$-even states.
@@
It has been demonstrated that $C$-even states can be produced in $e^{+}e^{-}$ collisions accompanied by photon emission. For instance, the $e^{+}e^{-}\to\gamma X(3872)$ gamma3872 and $\gamma\chi_{c1,2}(1P)$ gammachicj processes have been observed by the BESIII experiment. In this paper, a study of $e^{+}e^{-}\to\gamma D_{s}^{\pm}D_{s}^{*\mp}$ is presented to search for the aforementioned $C$-even structures. Throughout this paper, the charge conjugated mode is implied. This analysis utilizes a data sample collected at center-of-mass energy $\sqrt{s}=(4681.92\pm 0.08\pm 0.29)~{}\mathrm{MeV}$ with an integrated luminosity of $(1667.39\pm 0.21\pm 8.84)~{}\mathrm{pb}^{-1}$ BESIII:lum , where the first and second uncertainties are statistical and systematic, respectively. The $D_{s}^{*}$ meson is reconstructed using its radiative decay into the $\gamma D_{s}$ final state. The $D_{s}$ meson is reconstructed with its decay modes $D_{s}\to K^{+}K^{-}\pi$ and $K^{0}_{S}K$, with $K^{0}_{S}\to\pi^{+}\pi^{-}$. These two decay modes of the $D_{s}$ are combined into four decay chains (DCs):
@@
Simulated events produced with a geant4-based geant4 Monte Carlo (MC) package are used to determine detection efficiencies and to estimate backgrounds. This MC package includes the geometric description of the BESIII detector and the detector response. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc ref:kkmca ; ref:kkmcb . The generic MC samples include the production of open-charm processes, the ISR production of vector charmonium(like) states, and the continuum processes implemented in kkmc ref:kkmca ; ref:kkmcb . The known decay modes are modeled with evtgen ref:evtgen ; ref:evtgena using branching fractions taken from the Particle Data Group (PDG) PDG2018 , and the remaining unknown charmonium decays are modelled with lundcharm ref:lundcharma ; ref:lundcharmb . Final state radiation for charged final state particles is incorporated using the photos package ref:photos .
The signal processes $e^{+}e^{-}\to\gamma X,~{}X\to D_{s}^{\pm}D_{s}^{*\mp}$ and their subsequent decays are modeled with evtgen ref:evtgen ; ref:evtgena . The $D_{s}\to K^{+}K^{-}\pi$ decay is generated according to the amplitude models reported in Refs. ref:da1 ; ref:da2 . Two groups of samples are generated for the signal processes, Group-I and Group-II. Group-I is produced with the predicted masses and widths of the $X$ as listed in Table 1, while Group-II is with the masses ranging from $4080~{}\mathrm{MeV}/c^{2}$ to $4320~{}\mathrm{MeV}/c^{2}$ in steps of $40~{}\mathrm{MeV}/c^{2}$ and widths varying from a set of values: $5~{}\mathrm{MeV}$, $10~{}\mathrm{MeV}$, $20~{}\mathrm{MeV}$, $50~{}\mathrm{MeV}$, $80~{}\mathrm{MeV}$, and $100~{}\mathrm{MeV}$. The quantum numbers $J^{PC}$ of $X$ are listed in Table 1. In the signal MC simulations of $e^{+}e^{-}\to\gamma X$, $\gamma X$ are assumed to generate from $\psi(4660)$, and $\psi(4660)$ is parametrized using a Briet-Wigner function with parameters taken from the PDG PDG2018 .
Additional MC samples of $e^{+}e^{-}$$\to$$(\gamma_{\rm ISR})D_{s}^{(*)\pm}D_{s}^{*\mp}$, $(\gamma_{\rm ISR})D_{s}^{*}D_{s0}^{*}(2317)$, and $(\gamma_{\rm ISR})D_{s}^{(*)}D_{s1}(2460)$ events are generated to estimate the background contamination. They are modeled using helicity-amplitude models or phase space in evtgen ref:evtgen ; ref:evtgena . The production cross sections are taken from previous BESIII measurements ref:dsstardsstar ; ref:dsds2460 ; ref:dssds2460 .
@@
Each $K^{0}_{S}$ candidate is reconstructed from two oppositely charged tracks satisfying $|V_{z}|<20~{}\mathrm{cm}$ and $|\cos\theta|<0.93$. The two charged tracks are assigned as $\pi^{+}\pi^{-}$ without imposing PID criteria. They are constrained to originate from a common vertex and are required to have an invariant mass $(M_{\pi^{+}\pi^{-}})$ within the interval $(0.478,~{}0.518)$ GeV/$c^{2}$. The $\chi^{2}$ of the vertex fit is required to be less than $100$. A secondary vertex fit is performed to ensure that the $K^{0}_{S}$ momentum points back to the IP. The decay length of the $K^{0}_{S}$ candidate is required to be greater than twice the vertex resolution.
@@
Photon candidates are identified using isolated showers not associated with charged tracks in the EMC. The deposited energy of each shower must be greater than $25~{}\mathrm{MeV}$ in the barrel region ($|\cos\theta|<0.8$) and greater than $50~{}\mathrm{MeV}$ in the end cap regions ($0.920.86<|\cos\theta|<$ 0.92). The difference between the EMC time and the event collision time is required to be within $[0,~{}700]~{}\mathrm{ns}$ to suppress electronic noise and showers unrelated to the event.
@@
The selected $K^{\pm}$, $\pi^{\pm}$, and $K^{0}_{S}$ candidates in the event are combined to reconstruct $D_{s}\to K^{+}K^{-}\pi$ or $D_{s}\to K^{0}_{S}K$. The $D_{s}$ candidates are kept if the invariant masses of the $K^{+}K^{-}\pi$ or $K^{0}_{S}K$ systems, denoted as $M_{D_{s}}$, are within a ±$\pm 150~{}\mathrm{MeV}/c^{2}$ mass window around the nominal $D_{s}$ mass. The $D_{s}^{+}D_{s}^{-}$ pairs are selected requiring that each $K^{\pm}$, $\pi^{\pm}$, and $K^{0}_{S}$ candidate is used in one $D_{s}^{+}D_{s}^{-}$ pair at most once. If there is more than one pair with the same DC in an event, the one with $(M_{D_{s}^{+}}+M_{D_{s}^{-}})/2$ closest to the nominal mass of the $D_{s}$ meson PDG2018 is chosen.
@@
In addition to the $D_{s}^{+}D_{s}^{-}$ pair, each signal candidate is required to have at least two photons. A six-constraint (6C) kinematic fit is performed, where the four-momentum of the final state particles is constrained to that of the initial $e^{+}e^{-}$ system and the masses of the reconstructed $D_{s}^{+}$ and $D_{s}^{-}$ are constrained to their nominal masses. If more than two photons and (or) more than one $D_{s}^{+}D_{s}^{-}$ pair decaying in different DCs are found in an event, the combination with the minimum $\chi^{2}_{\rm 6C}$ is retained, and $\chi^{2}_{\rm 6C}$ is further required to be less than 200, where $\chi^{2}_{\rm 6C}$ is the $\chi^{2}$ from the 6C kinematic fit. To study the background contributions from non-$D_{s}$ processes, a four-constraint kinematic (4C) fit without the two $D_{s}$ mass constraints is applied, and the four-momenta of the final state particles are computed using the momenta determined by the 4C kinematic fit.
@@
To suppress background contributions from non-$D_{s}$ processes, a $D_{s}$ mass window of $|M_{D_{s}}-m_{D_{s}}|<21~{}\mathrm{MeV}/c^{2}$, which is about three times the mass resolution, is applied, where $m_{D_{s}}$ is the nominal $D_{s}$ mass. The two selected photon candidates and the two $D_{s}$ mesons can form four combinations of $\gamma D_{s}$. The one with invariant mass ($M_{\gamma D_{s}})$ closest to the nominal mass of the $D_{s}^{*}$ is retained for further analysis. The invariant mass $M^{\prime}_{\gamma D_{s}}=(M_{\gamma D_{s}}-M_{D_{s}}+m_{D_{s}})$ is required to be within the interval $(m_{D_{s}^{*}}\pm 27)~{}\mathrm{MeV}/c^{2}$, where $m_{D_{s}^{*}}$ is the nominal $D_{s}^{*\pm}$ mass taken from the PDG PDG2018 . The variable $M_{\gamma D_{s}}$ formed with the unused photon and $D_{s}$ candidates is required to be outside of this same mass window. The photon not from the $D_{s}^{*}$ decay is labeled $\gamma_{1}$ and its recoil mass spectrum, $M_{\gamma_{1}}^{\rm rec}$, is used to extract the signal yields. $M_{\gamma_{1}}^{\rm rec}$ is calculated by $\sqrt{(E_{\rm cms}-E_{\gamma_{1}})^{2}-(\vec{P}_{\rm cms}-\vec{P}_{\gamma_{1}})^{2}}$, where ($E_{\rm cms},\vec{P}_{\rm cms}$) and ($E_{\gamma_{1}},\vec{P}_{\gamma_{1}}$) are the four-momenta of the initial state BESIII:lum and $\gamma_{1}$, respectively. After imposing all the selection requirements, the dominant background processes are those that include one $D_{s}^{+}D_{s}^{-}$ pair in final states. The contributions from $e^{+}e^{-}$$\to$$(\gamma_{\rm ISR})D_{s}^{(*)\pm}D_{s}^{*\mp}$, $(\gamma_{\rm ISR})D_{s}^{*}D_{s0}^{*}(2317)$, and $(\gamma_{\rm ISR})D_{s}^{(*)}D_{s1}(2460)$ events account for $76.9\%$, $4.4\%$, and $10.9\%$ of the total backgrounds, respectively. The contributions from other background processes are less than $8\%$, and are found to be linearly distributed in the recoil mass spectrum of $\gamma_{1}$ based on the analysis of generic MC samples.
@@
The $X$ signal yields are extracted using an unbinned maximum likelihood fit to the $M_{\gamma_{1}}^{\rm rec}$ spectrum. The events from the four DCs are combined due to limited statistics. The fit probability density function (PDF) contains five components: signal ($S$), $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{\pm}D_{s}^{*\mp}$ background ($B_{1}$), $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{*\pm}D_{s}^{*\mp}$ background ($B_{2}$), $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{(*)}D_{sJ}$ background ($B_{3}$), and other remaining backgrounds ($B_{4}$). The function used in the fit is defined as
$f_{\rm sum}=N_{\rm sig}\cdot S+\sum_{i=1}^{4}N_{\mathrm{bkg}i}\cdot B_{i},$
where $N_{\rm sig}$ and $N_{\mathrm{bkg}(1,2,3,4)}$ represent the numbers of signal events and background events, respectively. $N_{\mathrm{bkg}(1,2,4)}$ are free parameters, and $N_{\mathrm{bkg}3}$ is fixed according to BESIII measurement in the fit, $N_{\mathrm{bkg}3}=8.8\pm 0.9$ ref:dsds2460 ; ref:dssds2460 , where the number of events and the uncertainty are from the cross section measurements.
The signal PDF is described by $S=BW(M_{\gamma_{1}}^{\rm rec};M,\Gamma)\otimes DG(m_{1},\sigma_{1},m_{2},\sigma_{2},f)\otimes G(\Delta m,\Delta\sigma)$, where $BW$ stands for the Breit-Wigner function with mass $M$ and width $\Gamma$ fixed, $DG$ is a double Gaussian function that describes the mass resolution and mass shift in the simulation, and $G$ is a Gaussian function that accounts for the mass resolution difference between data and MC simulation. The parameters $m_{1}(m_{2})$ and $\sigma_{1}(\sigma_{2})$ represent the mean and standard deviation of the first (second) Gaussian function in the double Gaussian model, while $f$ is the fraction of the first Gaussian component over the whole $DG$. To get the parameters of the double Gaussian function, the $M_{\gamma_{1}}^{\rm rec}-M_{\gamma_{1}}^{\rm truth}$ distribution from the signal MC samples is fitted, where $M_{\gamma_{1}}^{\rm truth}$ is the recoil mass spectrum of the $\gamma_{1}$ distribution at generator level. The parameters of the Gaussian function are obtained by using a control sample, $e^{+}e^{-}\to D_{s}^{\pm}D_{s}^{*\mp}$. The selection of the control sample is similar to that of the signal sample, except that only one photon is required instead of two. The $D_{s}$ and $D_{s}^{*}$ mass distributions from data are fitted with the line shape from MC samples convoluted with Gaussian functions. The parameters of the Gaussian function, representing the difference of the mass resolutions between data and MC simulation, are then used as the parameters of the Gaussian function in the extraction of the $X$ signal yield. The $B_{1(2)}$ line shape is taken from the $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{\pm}D_{s}^{*\mp}$ ($e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{*\pm}D_{s}^{*\mp}$) MC simulation, $B_{3}$ is from the sum of the $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{*}D_{s0}^{*}(2317)$ and $(\gamma_{\rm ISR})D_{s}^{(*)}D_{s1}(2460)$ MC simulations (weighted according to the product of the cross sections and detection efficiencies), and $B_{4}$ is modeled by a polynomial function.
@@
Fig. 1: Fits to the $M_{\gamma_{1}}^{\rm rec}$ spectrum with masses and widths fixed to those of the $\eta_{c2}(2D)$, $\chi_{c1}(3P)$, $\chi_{c2}(3P)$, $X(4080)$, $X(4217)$, and $X(4279)$ as listed in Table 1. The dots with error bars are the data distribution and the blue solid lines are the best fit results. The red arrows indicate the peak positions of those six $X$ states with masses fixed to the predictions listed in Table 1.
In addition, upper limits for the signal cross sections as a function of the $X$ masses in the range $[4.08,~{}4.32]~{}\mathrm{GeV}/c^{2}$ are also provided. From studies of simulation samples (Group-II), it is found the parameters $m_{2}-m_{1}$, $\sigma_{2}/\sigma_{1}$, and $f$ of the resolution function are very similar for various widths and $J^{PC}$. The parameters are found to be